# OphNet: A Large-Scale Video Benchmark for Ophthalmic Surgical Workflow Understanding

### v3

@@ Abstract
Surgical scene perception via videos are critical for advancing robotic surgery, telesurgery, and AI-assisted surgery, particularly in ophthalmology. However, the scarcity of diverse and richly annotated video datasets has hindered the development of intelligent systems for surgical workflow analysis. Existing datasets for surgical workflow analysis, which typically face challenges such as small scale, a lack of diversity in surgery and phase categories, and the absence of time-localized annotations, limit the requirements for action understanding and model generalization validation in complex and diverse real-world surgical scenarios. To address this gap, we introduce OphNet, a large-scale, expert-annotated video benchmark for ophthalmic surgical workflow understanding. OphNet features: 1) A diverse collection of 2,278 surgical videos spanning 66 types of cataract, glaucoma, and corneal surgeries, with detailed annotations for 102 unique surgical phases and 150 granular operations; 2) It offers sequential and hierarchical annotations for each surgery, phase, and operation, enabling comprehensive understanding and improved interpretability; 3) Moreover, OphNet provides time-localized annotations, facilitating temporal localization and prediction tasks within surgical workflows. With approximately 205 hours of surgical videos, OphNet is about 20 times larger than the largest existing surgical workflow analysis benchmark. Our dataset and code have been made available at: https://github.com/minghu0830/OphNet-benchmark.
@@ 1 Introduction
Automatic video surgical workflow understanding is a fundamental yet challenging problem for developing computer-assisted and robotic-assisted surgery, which can be divided into internal (e.g. laparoscopic and endoscopic [27, 46, 56]) and external (e.g. operating room and nursing procedure [40]) analysis. In addition to promoting the development of intelligent surgery, it also greatly benefits surgical documentation, education, and training [64, 11, 12]. Baret et al. [6] showed networks, like Inflated 3D ConvNet (I3D) [10] that utilize spatiotemporal convolutions, require a relatively extensive dataset for effective training. In their study, the model achieves an accuracy exceeding 80% when trained on 100 videos, with a progressive improvement as the sample size surpasses 700. However, the highly efficient and rapidly evolving deep learning technologies for surgical workflow analysis are currently limited by the following shortcomings in current video benchmarks: 1) Small scale: the majority of video datasets contain no more than 100 videos. For example, the CATARACTS [22] and CatRelDet [22] datasets contain only 50 and 21 surgical videos, respectively. These datasets are relatively small, insufficient for large-scale validation. 2) Limited categories of surgeries and phases: almost all ophthalmic surgical video datasets only include cataract surgery and do not further classify specific types of surgeries. Additionally, the number of phase categories is also limited, like CatRelDet [22] only contains 4 different phase labels, which is insufficient to meet the requirements for evaluation in real clinical environments. 3) Coarse-grained annotation: due to annotation costs, existing benchmarks often have coarse-grained action definitions. For example, adhesive injection may occur in two different phases: main incision and capsulorhexis, so it may be classified into different phase categories. Coarse-grained action definitions may lead to annotation bias. 4) Single time-boundary annotation: they only annotate designated phases in the videos, ignoring the continuity across different stages of ophthalmic surgery, as well as the hierarchical relationship between surgery, phase, and operation. Simpler datasets, such as LensID [21], are limited to binary classification tasks distinguishing lens implantation from other irrelevant phases. 5) Uniform domain: the videos are meticulously collected, and while this ensures video quality, the uniform style is not conducive to testing the model’s domain generalization ability.
@@
OphNet(Ours) 2,078 9,795 66 150 204.8h
Table 1: The statistics comparison among existing workflow analysis datasets and our OphNet. Compared to other datasets, OphNet focuses on more comprehensive coverage of various surgery, phase and operation categories, collects a large number of videos, totaling 204.8 hours, and also enables a variety of recognition, localization and prediction tasks. OphNet demonstrates considerable competitiveness in both its scale and the richness of its labels. Endo&Lap denotes the endoscopic and laparoscopic protocol, OphScope denotes the ophthalmic microscope protocol. We choose the latest version for comparison in cases where datasets have multiple supplementary updates. For instance, Cholec120 [44], Cholec80 [64], m2cai-workflow and LapChole [61] form one series, whereas CholecT50 [45], CholecT45 [45], and CholecT40 [43] comprise another series. We have excluded the following scenarios from our comparison: (1) non-open-source datasets such as Bypass170 [65], ESD [29], Yu’s [73], etc.; (2) a superset of multiple open-source or non-open-source datasets, like Cholec207 [6], etc.; (3) datasets employed for lesion, anatomy, and instrument classification and segmentation, such as SUN-SEG [26], CVC-ClinicDB [7], ROBUST-MIS [51], Mesejo’s [39], Cata7 [42], etc., anomaly detection such as PolypDiag [62] (from Hyper-Kvasir [9] and LDPolypVideo [37]), Kvasir-Capsule [58], etc., and other datasets not dedicated to workflow analysis. It’s worth mentioning that even in comparison with the above datasets, OphNet demonstrates considerable competitiveness in both its scale and the richness of its labels.
@@
    Fine-grained, sequential and hierarchical annotation: we have meticulously selected a subset of videos for annotation localization, with each video being marked for an average of 13 phases and 17 operations. Additionally, we provide exquisite annotations at the levels of surgery, phase, and operation, catering to the requirements for training specific challenge models. This annotation design aims to offer a multifaceted understanding of surgical protocols, accommodate the nuances of each distinct surgery, and enhance the usability and interpretability of our dataset.
@@ 2 Related Work
Surgical Workflow Understanding. Beyond its therapeutic advantages, minimally invasive surgery also provides the capability for operative video recording. These videos can be stored and later utilized for various purposes such as cognitive training, skill assessment, and surgical workflow analysis. Techniques derived from the broader field of video content analysis and representation are increasingly being incorporated into the surgical realm [35, 73, 6]. A typical surgical workflow can be defined by a sequence of tasks or events, including patient positioning, incision, dissection, and suturing. These events are influenced not only by the specific type of surgery but also by the individual surgeon’s proficiency and technique. Consequently, a comprehensive understanding of the surgical workflow necessitates a thorough examination of the temporal, spatial, and contextual facets of these tasks.
@@ 3 Dataset Construction
In this section, We describe the construction of the dataset involved meticulous data collection and preprocessing, leveraging YouTube as a primary source to circumvent privacy issues while ensuring a broad representation of ophthalmic surgeries. The selection criteria aimed at capturing a wide array of video qualities and styles, specifically targeting cataract, glaucoma, and corneal surgeries due to their prevalence in clinical settings. Efforts were made to refine the dataset by excluding videos of inadequate quality or those depicting non-human subjects. The annotation process was structured to reflect the complex nature of eye surgeries, incorporating hierarchical classification to account for the multiple conditions often treated within a single procedure. This was complemented by detailed localization annotations, delineating the distinct phases and techniques characteristic of ophthalmic operations, undertaken by a dedicated team of ophthalmologists to ensure accuracy and relevance.
@@ 3.2 Data Annotation
Hierarchical Classification Annotation. Different from the single-label classification in most natural video datasets [57, 17, 78, 24, 59], in ophthalmic surgery, physicians typically consider and implement multiple types of surgeries and phases based on the specific conditions and needs of the patient. This multifaceted approach stems from the complexity of the eye as an organ, where numerous diseases often coexist. Consequently, multiple ocular problems may need to be addressed in a single surgery. As shown in Fig. 2, a patient with cataracts may also have glaucoma, necessitating the simultaneous treatment of both conditions within one surgery. Hence, we initially classified all videos based on the primary surgical categories into three principal groups: cataract, glaucoma, and corneal surgeries. Subsequently, these were further allocated for the detailed annotation of primary and secondary surgery. It is important to note that there exists only a single type of primary surgery, whereas multiple secondary surgery may be present. The categorization and annotation were meticulously completed by a team comprised of 8 experienced ophthalmologists.
Localization Annotation. A single surgery is often multifaceted, involving a series of intricate phases, each requiring distinct techniques and instruments, and the transition between surgeries entails high precision and coordination. Routine cataract surgery involves several steps. It begins with the administration of anesthesia, followed by a small incision in the cornea. The surgeon then creates an opening in the lens capsule and uses ultrasonic vibrations to break up and remove the cataract. Afterward, an artificial intraocular lens (IOL) is inserted into the lens capsule, and the incision is sealed without stitches. We define the phases and operation for various surgeries based on the textbook Ophthalmic Surgery: Principles and Practice [60]. To ensure the quality of localization annotations, we assign each annotator videos of 2-6 different types of secondary surgeries. This process resulted in a minimum of three annotated action boundaries for each video. We also employed the complete linkage algorithm [13] to cluster and merge various temporal boundaries into stable boundaries that received multiple agreements. It’s worth noting that an individual video may have multiple separate instances of the same or different phases, thereby leading to multiple boundary definitions. A team of 15 ophthalmologists completes the localization annotation.
@@ 3.3 Dataset Statistics and Analysis
OphNet includes 2,078 surgical videos (204.8 hours), demonstrating 66 different types of ophthalmic surgeries: 13 types of cataract surgery, 14 types of glaucoma surgery, and 39 types of corneal surgery. There are 102 phases and 150 operations for recognition, detection and prediction tasks, summarized in Tab. 1. Over 77% of videos have high-definition resolutions of 1280 x 720 pixels or higher. To facilitate algorithm development and evaluation, we selected 523 videos for localization annotations. Additionally, we trim videos according to annotated action boundaries, resulting in 7,320 phase instances and 9,795 operation instances, totaling 51.2 hours. The average duration of trimmed videos is 32 seconds, while untrimmed videos average 337 seconds.
@@ 4 Experiments
In our study, we explore four potential tasks using the OphNet dataset: 1) surgery presence recognition, 2) phase and operation recognition, 3) phase localization, and 4) phase anticipation. To establish robust baselines, we employ state-of-the-art models known for their effectiveness in human action recognition, detection and anticipation. For each task, we provide a detailed problem formulation and evaluate the baseline models‘ performance. Our findings offer valuable insights into video understanding within sequences and fine granularity, contributing to the field’s knowledge of video tasks in medical contexts.
@@ 4.1 Surgery Presence, Phase and Operation Recognition
Task Description. Surgery presence recognition focuses on identifying various surgical types in untrimmed videos through weakly supervised methods. This requires the model to discern and capture distinct surgical action features within extensive video footage. In OphNet, every surgical procedure within each video is annotated. We identify the primary surgery based on the key objectives and durations of the surgeries. To streamline the process, our experiments are limited to recognizing the presence of these primary surgeries. Additionally, phase recognition segments the surgery into distinct phases using visual cues and movements, such as incision, lens removal, and implantation. Meanwhile, operation recognition involves identifying finer-grained surgical actions.
Setup. For the surgery presence recognition experiment, we simplified the process and only considered the presence of the primary surgery in untrimmed videos. The dataset was randomly partitioned to ensure a balanced representation of examples for each surgery category. Specifically, we allocated 70% of the data for training (1,449 surgical videos), 10% for validation (205 surgical videos), and 20% for testing (424 surgical videos). For the phase and operation recognition experiments, we maintained the same settings, with 70% of the data used for training (5,024 phase segments, 6,856 operations segments), 10% for validation (730 phase segments, 975 operations segments), and 20% for testing (1,566 phase segments, 1,964 operations segments). The input for the surgery presence recognition experiment is untrimmed videos, while for the phase and operation recognition experiments, the input consists of trimmed segments. In all classification experiments, we set up the analysis and comparison from four perspectives: cataract surgery, glaucoma surgery, corneal surgery, and all surgical videos.
@@
Baselins. We compare the performance of I3D [10], SlowFast [19], X3D [18], and MViT V2 [30] models on this task. These models are evaluated in two versions: 1) random initialization training and 2) pre-training with weights from Kinetics 400 [28], which is a human action recognition dataset. In addition, we also explored the classification performance of X-CLIP [41] and ViFi-CLIP [48], two CLIP [47]-based models. We also compared the effects of different numbers of input frames on the models’ performance, where the subscript "_16" represents an input of 16 frames, and "_32" represents an input of 32 frames.
@@
4.2 Phase Localization
@@
Baselins. We conducted the experiments using phase-level labels and used two baseline models, ActionFormer [76] and TriDet [55], with backbone networks configured as CSN [63], SwinViviT [33], and SlowFast [19]. Data split follows the setup of the primary surgery classification experiment.
@@
Figure 4: Attention map visualizations of ViFi-CLIP [48] on four examples from OphNet test set in the phase recognition task.
@@ 4.3 Phase Anticipation
Baselins. We evaluate our datasets with the baseline models such as I3D [10], SlowFast [19], and MViT V2 [30]. For each model, we also adopted two training approaches: random initialization training and using pre-trained weights from Kinetics 400 [28].
Results. We show the phase detection results in Tab. 5. The results demonstrate that the baseline models pretrained on Kinetics 400 [28] generally outperform their original counterparts in terms of Top-1 accuracy across different observation ratios for phase anticipation. Specifically, the modified MViT V2* model exhibits the highest improvement, achieving the best average Top-1 accuracy of 48.2%. Moreover, while all models show increased accuracy with higher observation ratios, indicating that more observed data contributes to better performance, the consistent improvement across all ratios for the enhanced models suggests effective modifications.

### v4

@@ Abstract
Surgical scene perception via videos is critical for advancing robotic surgery, telesurgery, and AI-assisted surgery, particularly in ophthalmology. However, the scarcity of diverse and richly annotated video datasets has hindered the development of intelligent systems for surgical workflow analysis. Existing datasets face challenges such as small scale, lack of diversity in surgery and phase categories, and absence of time-localized annotations. These limitations impede action understanding and model generalization validation in complex and diverse real-world surgical scenarios. To address this gap, we introduce OphNet, a large-scale, expert-annotated video benchmark for ophthalmic surgical workflow understanding. OphNet features: 1) A diverse collection of 2,278 surgical videos spanning 66 types of cataract, glaucoma, and corneal surgeries, with detailed annotations for 102 unique surgical phases and 150 fine-grained operations. 2) Sequential and hierarchical annotations for each surgery, phase, and operation, enabling comprehensive understanding and improved interpretability. 3) Time-localized annotations, facilitating temporal localization and prediction tasks within surgical workflows. With approximately 285 hours of surgical videos, OphNet is about 20 times larger than the largest existing surgical workflow analysis benchmark. Code and dataset are available at: https://minghu0830.github.io/OphNet-benchmark/.
@@ 1 Introduction
Automatic video surgical workflow understanding is a fundamental yet challenging problem for developing computer-assisted and robotic-assisted surgery, which can be divided into internal (e.g., laparoscopic and endoscopic [28, 47, 57]) and external (e.g., operating room and nursing procedure [41]) analysis. In addition to promoting the development of intelligent surgery, it also greatly benefits surgical documentation, education, and training [65, 11, 12]. Baret et al. [6] showed networks, like Inflated 3D ConvNet (I3D) [10] that utilize spatiotemporal convolutions, require a relatively extensive dataset for effective training. In their study, the model achieves an accuracy exceeding 80% when trained on 100 videos, with a progressive improvement as the sample size surpasses 700. However, the highly efficient and rapidly evolving deep learning technologies for surgical workflow analysis are currently limited by the following shortcomings in current video benchmarks: 1) Small-scale: the majority of video datasets contain no more than 100 videos. For example, the CATARACTS [23] and CatRelDet [23] datasets contain only 50 and 21 surgical videos, respectively. These datasets are relatively small, insufficient for large-scale validation. 2) Limited categories of surgeries and phases: almost all ophthalmic surgical video datasets only include cataract surgery and do not further classify specific types of surgeries. Additionally, the number of phase categories is also limited, like CatRelDet [23] only contains 4 different phase labels, which is insufficient to meet the requirements for evaluation in real clinical environments. 3) Coarse-grained annotation: due to annotation costs, existing benchmarks often have coarse-grained action definitions. For example, adhesive injection may occur in two different phases: main incision and capsulorhexis, so it may be classified into different phase categories. Coarse-grained action definitions may lead to annotation bias. 4) Single time-boundary annotation: they only annotate designated phases in the videos, ignoring the continuity across different stages of ophthalmic surgery, as well as the hierarchical relationship between surgery, phase, and operation. Simpler datasets, such as LensID [22], are limited to binary classification tasks distinguishing lens implantation from other irrelevant phases. 5) Uniform domain: the videos are meticulously collected, and while this ensures video quality, the uniform style is not conducive to testing the model’s domain generalization ability.
@@
Cataract-1K [21] 1,000 931 1 12 118.7h
OphNet(Ours) 2,278 9,795 66 150 284.8h
Table 1: The statistics comparison among existing workflow analysis datasets and our OphNet. Compared to other datasets, OphNet focuses on more comprehensive coverage of various surgery, phase and operation categories, collects a large number of videos, totaling 284.8 hours, and also enables a variety of recognition, localization and prediction tasks. OphNet demonstrates considerable competitiveness in both its scale and the richness of its labels. For instance, Cholec120 [45], Cholec80 [65], m2cai-workflow and LapChole [62] form one series, whereas CholecT50 [46], CholecT45 [46], and CholecT40 [44] comprise another series. We have excluded the following scenarios from our comparison: (1) non-open-source datasets such as Bypass170 [66], ESD [30], Yu’s [74], etc.; (2) a superset of multiple open-source or non-open-source datasets, like Cholec207 [6], etc.; (3) datasets employed for lesion, anatomy, and instrument classification and segmentation, such as SUN-SEG [27], CVC-ClinicDB [7], ROBUST-MIS [52], Mesejo’s [40], Cata7 [43], etc., anomaly detection such as PolypDiag [63] (from Hyper-Kvasir [9] and LDPolypVideo [38]), Kvasir-Capsule [59], etc., and other datasets not dedicated to workflow analysis. It’s worth mentioning that even in comparison with the above datasets, OphNet demonstrates considerable competitiveness in both its scale and the richness of its labels. Endo&Lap denotes the endoscopic and laparoscopic protocol, OphScope denotes the ophthalmic microscope protocol. We choose the latest version for comparison in cases where datasets have multiple supplementary updates.
@@
    Fine-grained, sequential and hierarchical annotation: we have meticulously selected a subset of videos for annotation localization, with each video being annotated for an average of 22 operations. Additionally, we provide exquisite annotations at the levels of surgery, phase, and operation, catering to the requirements for training specific challenge models. This annotation design aims to offer a multifaceted understanding of surgical protocols, accommodate the nuances of each distinct surgery, and enhance the usability and interpretability of our dataset.
@@ 2 Related Work
Surgical Workflow Understanding. Beyond its therapeutic advantages, minimally invasive surgery also provides the capability for operative video recording. These videos can be stored and later utilized for various purposes such as cognitive training, skill assessment, and surgical workflow analysis [76, 77]. Techniques derived from the broader field of video content analysis and representation are increasingly being incorporated into the surgical realm [36, 74, 6]. A typical surgical workflow can be defined by a sequence of tasks or events, including patient positioning, incision, dissection, and suturing. These events are influenced not only by the specific type of surgery but also by the individual surgeon’s proficiency and technique. Consequently, a comprehensive understanding of the surgical workflow necessitates a thorough examination of the temporal, spatial, and contextual facets of these tasks.
@@ 3 Dataset Construction
In this section, we detail the construction of our dataset, which involved meticulous data collection and preprocessing. YouTube was leveraged as a primary source to circumvent privacy issues while ensuring a broad representation of ophthalmic surgeries. Our selection criteria aimed to capture a wide range of video qualities and styles, specifically targeting cataract, glaucoma, and corneal surgeries due to their prevalence in clinical settings. We refined the dataset by excluding videos of inadequate quality or those depicting non-human subjects. The annotation process was designed to reflect the complex nature of eye surgeries, incorporating hierarchical classification to account for multiple conditions often treated within a single procedure. This was complemented by detailed localization annotations, delineating the distinct phases and techniques characteristic of ophthalmic operations, all undertaken by a dedicated team of ophthalmologists to ensure accuracy and relevance.
@@ 3.2 Data Annotation
Classification Annotation. Different from the single-label classification in most natural video datasets [58, 17, 81, 25, 60], in ophthalmic surgery, physicians typically consider and implement multiple types of surgeries and phases based on the specific conditions and needs of the patient. This multifaceted approach stems from the complexity of the eye as an organ, where numerous diseases often coexist. Consequently, multiple ocular problems may need to be addressed in a single surgery. As shown in Fig. 1, a patient with cataracts may also have glaucoma, necessitating the simultaneous treatment of both conditions within one surgery. Hence, we initially classified all videos based on the primary surgical categories into three principal groups: cataract, glaucoma, and corneal surgeries. Subsequently, these were further allocated for the detailed annotation of primary and secondary surgery. It is important to note that there exists only a single type of primary surgery, whereas multiple secondary surgery may be present. The categorization and annotation were meticulously completed by a team comprised of 8 experienced ophthalmologists.
Hierarchical Localization Annotation. A single surgery is often multifaceted, involving a series of intricate phases, each requiring distinct techniques and instruments, and the transition between surgeries entails high precision and coordination. Routine cataract surgery involves several steps. It begins with the administration of anesthesia, followed by a small incision in the cornea. The surgeon then creates an opening in the lens capsule and uses ultrasonic vibrations to break up and remove the cataract. Afterward, an artificial intraocular lens (IOL) is inserted into the lens capsule, and the incision is sealed without stitches. We define the phases and operation for various surgeries based on the textbook Ophthalmic Surgery: Principles and Practice [61]. To ensure the quality of localization annotations, we assign each annotator videos of 2 different types of primary surgeries. For each action time-boundary annotation, we annotate at three different levels of granularity: surgery, phase, and operation. We also employed the complete linkage algorithm [13] to cluster and merge various temporal boundaries into stable boundaries that received multiple agreements. It’s worth noting that an individual video may have multiple separate instances of the same or different phases, thereby leading to multiple boundary definitions. A team of 15 ophthalmologists completes the localization annotation.
@@ 3.3 Dataset Statistics and Analysis
OphNet includes 2,278 surgical videos (284.8 hours), demonstrating 66 different types of ophthalmic surgeries: 13 types of cataract surgery, 14 types of glaucoma surgery, and 39 types of corneal surgery. There are 102 phases and 150 operations for recognition, detection and prediction tasks, summarized in Tab. 1. Over 77% of videos have high-definition resolutions of 1280 $\times$ 720 pixels or higher. To facilitate algorithm development and evaluation, we selected 523 videos for localization annotations. Additionally, we trim videos according to annotated action boundaries, resulting in 7,320 phase instances and 9,795 operation instances, totaling 51.2 hours. The average duration of trimmed videos is 32 seconds, while untrimmed videos average 337 seconds.
@@ 4 Experiments
In our study, we explore four potential tasks using the OphNet dataset: 1) primary surgery presence recognition, 2) phase and operation recognition, 3) phase localization, and 4) phase anticipation. To establish robust baselines, we employ state-of-the-art models known for their effectiveness in human action recognition, detection and anticipation. For each task, we provide a detailed problem formulation and evaluate the baseline models‘ performance. Our findings offer valuable insights into video understanding within sequences and fine granularity, contributing to the field’s knowledge of video tasks in medical contexts.
@@ 4.1 Surgery Presence, Phase and Operation Recognition
Task Description. Surgery presence recognition focuses on identifying various surgical types in untrimmed videos through weakly supervised methods. This requires the model to discern and capture distinct surgical action features within extensive video footage. In OphNet, all types of surgeries occurring in each video are annotated, but only a subset of the videos have time-boundary annotations, as detailed in Sec. 3.2. We identify the primary surgery based on the key objectives and durations of the surgeries. To streamline the process, our experiments are limited to recognizing the presence of these primary surgeries. Additionally, phase recognition segments the surgery into distinct phases using visual cues and movements, such as incision, lens removal, and implantation. Operation recognition involves identifying finer-grained surgical actions.
@@
Baselins. We compare the performance of I3D [10], SlowFast [19], X3D [18], and MViT V2 [31] models on this task. These models are evaluated in two versions: 1) random initialization training and 2) pre-training with weights from Kinetics 400 [29], which is a human action recognition dataset. In addition, we also explored the classification performance of X-CLIP [42] and ViFi-CLIP [49], two CLIP [48]-based models. We also compared the effects of different numbers of input frames on the models’ performance, where the subscript _16 represents an input of 16 frames, and _32 represents an input of 32 frames.
Setup. The dataset was randomly partitioned to ensure a balanced representation of examples for each surgery category. Specifically, we allocated 70% of the data for training (1,449 surgical videos), 10% for validation (205 surgical videos), and 20% for testing (424 surgical videos). For the phase and operation recognition experiments, we followed the same settings, with 70% of the data used for training (5,024 phase segments, 6,856 operations segments), 10% for validation (730 phase segments, 975 operations segments), and 20% for testing (1,566 phase segments, 1,964 operations segments). The input for the surgery presence recognition experiment is untrimmed videos, while for the phase and operation recognition experiments, the input consists of trimmed segments. In all classification experiments, we set up the analysis and comparison from four perspectives: cataract surgery, glaucoma surgery, corneal surgery, and all surgical videos.
@@ 4.2 Phase Localization
Baselines. We conducted the experiments using phase-level labels and used two baseline models, ActionFormer [79] and TriDet [56], with backbone networks configured as CSN [64], SwinViviT [34], and SlowFast [19]. Data split follows the setup of the primary surgery classification experiment.
@@
Figure 4: Attention map visualizations of ViFi-CLIP [49] on four examples from OphNet’s test set in the phase recognition task.
@@ 4.3 Phase Anticipation
Baselines. We evaluate our datasets with the baseline models such as I3D [10], SlowFast [19], and MViT V2 [31]. For each model, we also adopted two training approaches: random initialization training and using pre-trained weights from Kinetics 400 [29].
Results. The phase detection results are illustrated in Tab. 5. The results demonstrate that the baseline models pretrained on Kinetics 400 [29] generally outperform their original counterparts in terms of Top-1 accuracy across different observation ratios for phase anticipation. Specifically, the modified MViT V2* model exhibits the highest improvement, achieving the best average Top-1 accuracy of 48.2%. Moreover, while all models show increased accuracy with higher observation ratios, indicating that more observed data contributes to better performance, the consistent improvement across all ratios for the enhanced models suggests effective modifications.
[…]
Figure 5: Phase localization visualization of TriDet [56]. GT represents the ground truth visualization for phases, while Loc. visualizes the model’s highest confidence phase category and the time-boundary results. Blank segments denote invalid segments or operation gaps.